find all zeros of the function. h(x)=x^3+4x^2+x-6
Answer by ewatrrr(24785)


Hi
finding all zeros of the function.
h(x)=x^3+4x^2+x-6
x^3+4x^2+x-6 = 0 |If x=1 then zero (1 + 4 + 1 - 6 = 0 )
(x -1)(x+3)(x+2) =0 |Dividing x^3+4x^2+x-6 by (x-1)
(x-1)=0 x= 1
(x+3)=0 x= -3
(x+2) =0 x = -2